# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-tuning

### v1

@@ B.4 Evaluation Setup Details
B.5 Other Safety Evaluation
In Table A1, we present additional safety evaluations of different safety fine-tuning approaches using LLaVA-v1.5-7B on two additional safety benchmarks: MM-SafetyBench and FigStep. The results are consistent with our earlier findings on VLGuard and SPA-VL: The one-word attack leads to a significant increase in ASR on both MM-SafetyBench and FigStep. This reveals that the spurious correlations learned during conventional safety fine-tuning persist across different evaluation settings. Our proposed MU-based methods demonstrate improved robustness, thereby alleviating this vulnerability.
Table A1: Safety evaluation on MM-Safetybench and FigStep. ASR is reported for unsafe inputs before and after the “What”-initialized 3-shot attack. The setup and format follow Table 1.
@@ B.5 Other Safety Evaluation
Table A2: Evaluating the safety and over-prudence under various visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups are consistent with Table 1.
Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence (RR, $\downarrow$)
@@
B.6 Examples of Input Saliency
B.7 Spurious Correlations with Visual Variations
To further evaluate the persistence of spurious correlations between text queries and safety labels and the robustness of our MU-based safety fine-tuning approaches, we present in Table A2 the safety performance of fine-tuned models when subjected to various visual perturbations. The results reveal that the SFT-based safety models consistently exhibit a significant increase in both ASR and RR even under visual variations. This finding indicates that the spurious correlations, and consequently the safety mirage phenomenon remains stable. In contrast, the performance of our MU-based safety aligned models remains robust across these perturbations.

### v2

@@ D.3 Evaluation Setup Details
D.4 Balance between Safety and Utility
Table A1: Overall results evaluating safety (ASR, $\downarrow$), over-prudence ($R$, $\downarrow$), and utility (Acc., $\uparrow$) on VQA benchmarks. Metrics are reported before and after one-word attacks/modifications. We compare conventional fine-tuning baselines, unlearning-based methods (NPO, RMU), and their coreset variants (50%, 25%, 10%). The other settings stay consistent with Table  1.
VLGuard SPA-VL VLGuard SPA-VL VQAv2 TextVQA ScienceQA VizWiz
Before After Before After Before After Before After
      + 50% NPO 2.94% 12.46% 19.02% 25.10% 2.69% 12.23% 17.45% 18.02% 77.41% 57.82% 68.20% 50.55%
      + 25% NPO 3.65% 14.25% 20.12% 26.48% 2.51% 11.33% 17.88% 18.67% 77.60% 58.10% 68.45% 50.92%
      + 10% NPO 5.17% 18.77% 22.01% 28.92% 1.97% 7.72% 18.32% 19.15% 78.10% 57.99% 68.72% 51.34%
      + 50% RMU 1.74% 10.64% 18.12% 23.05% 1.61% 8.01% 18.56% 19.70% 77.23% 57.98% 67.84% 50.34%
      + 25% RMU 3.22% 15.43% 19.06% 24.12% 1.61% 4.28% 18.92% 20.01% 78.01% 57.24% 68.12% 50.78%
      + 10% RMU 7.51% 22.12% 21.35% 27.25% 0.72% 1.97% 19.48% 20.73% 78.32% 58.97% 68.65% 51.42%
To better balance safety and utility in VLMs, we further adopt the idea of coreset unlearning (Pal et al., 2025). Instead of unlearning all safety data, we fine-tune on a random-sampled subset (e.g., 50%, 25%, or 10%) of the safety dataset. As shown in Table A1, reducing the coreset size leads to slightly higher ASR and RR (weaker safety), but improves accuracy on downstream tasks (higher utility). This also reflects a natural trade-off: smaller coresets sacrifice some safety robustness while preserving more general task capability. Importantly, even with only 10% of the safety data, unlearning-based models still achieve substantially better safety–utility trade-offs than conventional fine-tuning baselines, confirming the flexibility of coreset unlearning for practical deployment
Figure A4: Safety evaluation on MM-SafetyBench and FigStep. ASR is reported before and after the “What”-based 3-shot attack. Setup follows Table 1.
@@ D.4 Balance between Safety and Utility
D.5 Additional Safety Evaluation
In Table A4, we provide additional safety evaluations of different fine-tuning approaches for LLaVA-v1.5-7B on two further safety benchmarks: MM-SafetyBench and FigStep. The results align closely with our earlier findings on VLGuard and SPA-VL. In both new benchmarks, the one-word attack substantially increases the attack success rate (ASR), indicating that spurious correlations between surface-level textual cues and safety labels, induced during conventional safety fine-tuning, persist across diverse evaluation settings. These correlations cause models to become vulnerable to simple adversarial word insertions, undermining their safety guarantees. By contrast, our MU-based approaches (NPO and RMU) consistently achieve lower post-attack ASR, demonstrating improved robustness and reduced reliance on spurious features. This highlights that MU not only mitigates overfitting to dataset-specific biases but also generalizes better across benchmarks, thereby offering a more reliable defense against spurious correlation–driven vulnerabilities.
D.6 Examples of Input Saliency
Figure A7: Evaluating safety (ASR, $\downarrow$) and over-prudence ($R$, $\downarrow$) under visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups follow Table 1.
Models Safety (ASR) Over-Prudence (RR)
@@ D.6 Examples of Input Saliency
D.7 Spurious Correlations with Visual Variations
To further examine whether the identified spurious correlations between textual queries and safety labels persist under different visual conditions, we evaluate fine-tuned models on VLGuard with a range of common visual perturbations, including Gaussian Noise, Gaussian Blur, and Color Jitter. The results are summarized in Table,A7. We make several observations. First, conventional SFT-based safety models (Mixed-SFT and Posthoc-SFT) exhibit a consistent pattern of vulnerability: both ASR and RR remain high across all perturbations, showing that the learned spurious correlations are insensitive to low-level visual variations. This suggests that the safety mirage is not simply a dataset artifact tied to specific image statistics, but a systematic bias reinforced during fine-tuning. Second, MU-based approaches (NPO and RMU) maintain stable and low ASR and RR even under perturbations, confirming that unlearning not only alleviates spurious correlations in clean settings but also provides robustness against visual shifts. Third, comparing perturbation types, we find that color-based distortions (Color Jitter) lead to slightly larger fluctuations than Gaussian Noise or Blur, though the overall trend remains the same. This indicates that safety mirage in SFT models is strongly text-driven and less sensitive to visual details, while MU-based approaches decouple safety predictions from such spurious text–label shortcuts. Taken together, these findings highlight that spurious correlations embedded during safety fine-tuning are persistent and modality-invariant: they survive visual perturbations that alter low-level image statistics. In contrast, unlearning-based models generalize more reliably across such shifts, further validating their robustness in multimodal safety alignment.
Appendix E The Use of LLMs
This work makes limited use of LLMs. Specifically, LLMs were employed exclusively for grammar correction and stylistic polishing of the manuscript. They were not involved in research ideation, experimental design, data analysis, or the generation of any scientific content. All substantive contributions to the paper are solely attributable to the author.
Appendix F Limitations and broad Impact
F.1 Limitations
Our analysis has so far been limited to models at the 7B and 13B scales. Extending the study to larger foundation models remains an important next step, as scaling may introduce new behaviors or amplify new challenges. In addition, our experiments have primarily relied on widely used safety benchmarks. While these datasets provide systematic evaluation settings, they may not fully capture the diversity and complexity of real-world applications. Future work should therefore investigate unlearning performance in broader, domain-specific scenarios and corresponding downstream tasks.
F.2 Broad Impact
This work sheds light on the reliability of current VLM safety fine-tuning, uncovering the risk of a “safety mirage” where perceived robustness is illusory. By identifying spurious correlations as a root cause and demonstrating how machine unlearning can alleviate them, our study contributes to building safer and more trustworthy multimodal systems. In practice, stronger defenses against jailbreaking and over-prudence can reduce the risk of harmful or overly restrictive outputs, improving user trust in deployed AI systems. At the same time, unlearning techniques must be carefully applied, as overly aggressive removal of knowledge could impair model utility or be misused to deliberately suppress content. We encourage future research to balance safety, fairness, and transparency when deploying MU-based safety interventions in real-world applications.